OpenStudy (anonymous):

Which is always a correct conclusion about the quantities in the function y=x+8? A. The variable x is always 8 more than y. B. The variable y is always greater than x. C. As the value of x increases,the value of y decreases. D. When the value of x is negative, the value of y is also negative.

OpenStudy (anonymous):

I think the answer is c

OpenStudy (anonymous):

I would first eliminate the obvious two C and D add 1 to 8 and y increases make x= -1 and y is positive

OpenStudy (anonymous):

not C

jimthompson5910 (jim_thompson5910):

What's the slope of y=x+8 ?

OpenStudy (anonymous):

X is the slope

jimthompson5910 (jim_thompson5910):

go back to y = mx+b m is the slope

jimthompson5910 (jim_thompson5910):

example y = 3x - 10 is in the form y = mx+b where m = 3 b = -10 so the slope of y = 3x - 10 is m = 3

jimthompson5910 (jim_thompson5910):

you can write y = x+8 as y =1x + 8

OpenStudy (anonymous):

I still don't get it I've never been good in this subject

jimthompson5910 (jim_thompson5910):

basically the slope is just the number in front of x (the coefficient of x)

jimthompson5910 (jim_thompson5910):

it will visually tell you if the line is going up (increasing) or going down (decreasing) as you move from left to right

OpenStudy (anonymous):

So it's 1x the slope

jimthompson5910 (jim_thompson5910):

just 1, not 1x

jimthompson5910 (jim_thompson5910):

1 is the slope that's positive so this line is increasing

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

The answer is A

jimthompson5910 (jim_thompson5910):

let's say x = 1 what is y when x = 1 ?

OpenStudy (anonymous):

Idk

jimthompson5910 (jim_thompson5910):

use y =x+8 to find out

jimthompson5910 (jim_thompson5910):

plug in x = 1 and evaluate

OpenStudy (anonymous):

I don't have a calculator

jimthompson5910 (jim_thompson5910):

all you're doing is replacing and adding

jimthompson5910 (jim_thompson5910):

y = x+8 y = 1+8 y = ?

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

x = 1, y = 9 is "The variable x is always 8 more than y" true?

OpenStudy (anonymous):

True

jimthompson5910 (jim_thompson5910):

x = 1 is 8 more than y = 9 ?

OpenStudy (anonymous):

The answer is a because it 8 more than y

jimthompson5910 (jim_thompson5910):

1 is not 8 more than 9 if it was stated the other way around, then it would be true

OpenStudy (anonymous):

It's b